OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, –1).

OpenStudy (angela210793):

perpendicular means tht the slope will be =-1/(slope of the line ur given)==1/-3=1/3 equation whn u have the slope and a point is y-yo=m(x-x0) y+1=1/3(x-6) y=x/3 -3